# Symmetry-Projected Spin-AGP Methods Applied to Spin Systems

### v1

@@ I Introduction
Within the condensed matter physics community, spin model Hamiltonians have long served as a primary framework to study frustrated quantum magnetism [1, 2]. With a few exceptions [3], spin lattice models beyond one dimension are not exactly solvable, requiring the use of approximate numerical approaches instead. Analogously to the Hartree-Fock method used in the electronic structure theory of fermionic systems, spin wave functions based on a single spin configuration are insufficient for strongly correlated systems. Recent research has shown that the spin antisymmetrized geminal power (spin-AGP) method, which extends the AGP [4, 5] concept from systems of paired fermions to spin systems, may provide a good reference state for spin systems [6]. However, it is notable that, despite yielding some good results for one-dimensional (1D) systems, spin-AGP and correlated methods based on it are less accurate for lattices with more complicated geometry.
@@ II.2 Space group symmetry projection
$P_{LM}$ and $P_{PG}$ are the lattice momentum symmetry projection operator and point group symmetry projection operator, respectively. $T_{\mathbf{r}}$ is the translational operator with $\mathbf{q}$ being the lattice momentum projected onto. $R_{l}$ is a symmetry operator of the corresponding point group and $\chi_{k}(l)$ is the character of the operator $R_{l}$ in irreducible representation $k$ of the lattice point group.
Note that space group operators in practice permute sites of the lattice. Therefore, space group operators acting on AGP produce another AGP whose geminal coefficients eta are permuted:
@@
For example, if the symmetry operator $G_{T_{1}}$ is the translation operator that translates site $i$ into $i+1$, and $\{\eta_{1},\eta_{2},\cdots,\eta_{M}\}$ is the geminal coefficient of the original AGP, we have $|\mathrm{AGP}(\eta_{G})\rangle=G_{T_{1}}|\mathrm{AGP}(\eta)\rangle=G_{T_{1}}|\eta_{1},\eta_{2},\cdots,\eta_{M}\rangle=|\eta_{2},\eta_{3},\cdots,\eta_{1}\rangle$.
@@
We note that the bra state, instead of using the projected AGP, is simplified to the original AGP without permutation. Only the ket state is the projected AGP, represented by a linear combination of non-orthogonal AGPs (LC-AGP). Since the number of space group symmetry operators, and hence the number of AGPs required for this LC-AGP, is typically $\mathcal{O}(M)$, where $M$ is the number of spin sites, the cost of space group symmetry-projected AGP is one factor of $M$ larger than that of the mean-field optimized AGP, making it $\mathcal{O}(M^{4})$. This is in contrast to a more general LC-AGP–style wave function, for which one needs to evaluate Hamiltonian overlaps between $\mathcal{O}(M^{2})$ pairs of AGPs, at an $\mathcal{O}(M^{5})$ computational cost.
@@ II.3 Complex Conjugate, Spin-Flip, and Time-Reversal Symmetry
where the complex conjugation operator $K$ is anti-unitary operator and satisfies
@@
Because this set is closed, time-reversal symmetry is automatically projected once we apply the complex conjguation and spin flip projectors.
@@ III.1 One-dimensional XXZ model
Figure 1: Energy of the 12-site 1D XXZ model with periodic boundary conditions by space group projected AGP methods. SG-AGP with real $\eta$ is denoted as real-SG-AGP, while complex-LM-AGP and complex-SG-AGP represent LM-AGP and SG-AGP with complex $\eta$, respectively. These methods are compared with mean-field optimized AGP.
The space group projected AGP is optimized with both real and complex geminal coefficient $\eta$, the energies of which are shown in Fig. 1. The result of lattice momentum projected AGP (LM-AGP), which projects lattice momentum symmetry but not point group symmetry, with complex $\eta$ is also included. The results are compared with mean-field optimized AGP. It can be seen that with complex $\eta$, the energy result is much superior to that of real $\eta$ at the cost of complex conjugation symmetry breaking. The error of complex SG-AGP methods are shown in Fig.2 and are compared with J2 coupled cluster (J2CC) [23], a correlated method based on a single AGP state. We note that the symmetry projection makes a huge difference in the energy results, and the complex-SG-AGP, as a symmetry-projected mean-field method, already produces a result comparable to correlated methods.
@@
Of course the energy is not the only quantity of interest. Correlation functions, in particular, are important and can help us understand the nature of magnetism. A single AGP is generally inadequate here, as it provides exact correlation functions at $\Delta=-1$ but predicts that the correlation function is independent of Delta for $|\Delta|\lesssim 1$.
The $S_{z}S_{z}$ spin-spin correlation is defined as
@@
where $i$ and $j$ are indices for sites. Results are shown in Fig. 3. It can be seen that the spin-spin correlation given by SG-AGP is quite close to exact, and as $\Delta$ goes from $-1$ to $1$, the system approaches a Néel structure.
Figure 3: $S_{z}$-$S_{z}$ correlation functions for the 12-site XXZ model with PBC, compared those from SG-AGP with exact results. The lines are for exact results, and the circles are SG-AGP. Different colors represent different $\Delta$ values.
@@ III.2 Two-dimensional $\mathrm{J_{1}-J_{2}}$ model
where $\langle pq\rangle$ and $\langle\langle pq\rangle\rangle$ denote sites $p$ and $q$ being nearest neighbors and next-nearest neighbors respectively. The space group projected AGP methods are tested on the square and triangular lattices, both of which show rich physics and have spin-liquid [24] phases for intermediate $J_{2}$.
@@ III.2.1 Square lattice
Figure 5: Energy of the $4\times 4$ square $\mathrm{J_{1}-J_{2}}$ model by space group projected AGP methods compared with mean-field optimized AGP and J2CC.
The energy of the $4\times 4$ square $\mathrm{J_{1}-J_{2}}$ model given by symmetry-projected AGP methods is shown in Fig. 5. We remind our readers that these energies are optimized with complex $\eta$ via the variation after projection approach. In addition to SG-AGP, we include the results from the combination of space group projection with complex conjugate and spin-flip projection, referred to as SGKF-AGP. The results are compared with J2CC, which, despite the discontinuity in the highly frustrated region around $\Delta\sim 0.5$, provides some of our best results for correlation atop a single AGP. We see that with the symmetry projections, SG-AGP and SGKF-AGP can already surpass the performance of correlated methods based on a single AGP.
@@
We observe that when $\mathrm{J_{2}}=0$, there is a negative spin-spin correlation for the nearest neighbors, while the second-nearest neighbors exhibit a positive correlation. This is indicative of the checkerboard Néel phase. For $\mathrm{J_{2}}=1$, the second-nearest-neighbor correlation becomes negative while that of nearest-neighbor tends toward 00, corresponding to the striped order. The projected AGP spin-spin correlation functions coincide well with the exact one for large and small $\mathrm{J_{2}}$. In the spin liquid phase with intermediate $\mathrm{J_{2}}$ values, the spin-spin correlation is qualitatively correct but shows some deviations for certain site distances.
Figure 6: $S_{z}$-$S_{z}$ correlation from various projected methods in the $4\times 4$ square $\mathrm{J_{1}-J_{2}}$ model with PBC, compared with exact results. Lines are for the exact result, cross for SG-AGP, and circle for SGKF-AGP. Different colors represent different $\mathrm{J_{2}}$ values. The horizontal axis $j$ is arranged according to the distance between two spins where $j$ can be regarded as $j$th-nearest-neighbor of site 00, as shown in the top right of the figure, where site 0 is colored black and 1—5 are the $j$th-nearest-neighbors, colored differently according to $j$.
@@ III.2.2 Triangular lattice
A typical 2D triangular lattice has a $D_{6}$ point group symmetry with 12 symmetry operations. For the same point symmetry to be obeyed, a finite-size lattice must be chosen to take the shape of either a rhombus or hexagon, as shown in Fig. 4. Lattice momentum symmetry involves translation along the 2 lattice vectors. The $D_{6}$ point group symmetry and the translation symmetry form the $p6m$ space group for the 2D triangular lattice lattice and the SG-AGP for an $M$-site lattice is a linear combination of $12M$ AGP states.
@@
It is worth mentioning that in order to correctly capture the 120° phase, the lattice must have a multiple of 3 sites along each translation vector. We will test our methods on the 12-site and 36-site lattices, which satisfy this restriction, and we will also check its behavior on the 16-site lattice breaks this restriction as shown in the bottom left of Fig. 4: if different colors represent spins spaced by 120°, adjacent sites across the periodic boundary will showcase the same color, thereby breaching the 120° phase.
@@
Figure 7: Energy of the 12-site triangular $\mathrm{J_{1}-J_{2}}$ model by space group projected AGP methods compared with mean-field optimized AGP.
@@
The 12-site lattice has a hexagonal shape, as shown in the top right of Fig. 4. The energy of space group projected AGP methods is shown in Fig. 7. Although a single mean-field optimized AGP provides only a qualitative insight, including additional symmetry projections brings the energy closer to the exact value, with the SG-AGP exhibiting an error of $10^{-2}$. Moreover, the exact energy is obtained when the complex conjugate projection is included, which might be attributed to the limited size of the system. The spin-spin correlation of SG-AGP for this lattice is illustrated in Fig. 8, demonstrating that SG-AGP accurately captures the correlation.
@@
Figure 8: Spin-Spin Correlation of SG-AGP for the 12-site triangular $\mathrm{J_{1}-J_{2}}$ model compared with exact results. The lines are for the exact results, and the circles is for SG-AGP. Different colors represent different $\mathrm{J_{2}}$ values. The horizontal axis $j$ is arranged according to the distance between two spins where $j$ can be regarded as $j$th-nearest-neighbor of site 00, as shown in the top right of the figure, where site 0 is colored black and 1—3 are the $j$th-nearest-neighbors, colored differently according to $j$.
@@
The space group projected AGP methods are also tested on the 16-site ($4\times 4$ rhombus) and 36-site ($6\times 6$ rhombus) triangular lattice shown in Fig. 9 and Fig. 10 respectively. In the 16-site lattice, where the lattice size is not a multiple of 3, due to the finite-size effect mentioned earlier, the energy curves of both the symmetry-projected AGP methods and exact solution deviate from those observed in the 12-site and 36-site lattice models. Despite the absence of the 120° phase, the exact ground state of this lattice continues to exhibit the same space group symmetry as other system sizes. The space group projected AGP methods can accurately capture the energy, particularly SGKF-AGP, which has an error of approximately $10^{-1}$.
@@
However, an important limitation of projected AGP methods (and of AGP itself) is that they are not size extensive. In the limit of large lattices, the energy per site reduces to that of the spin-BCS. Figure 10 shows results for the 36-site rhombus lattice. We can see that SG-AGP is significantly less accurate here than it is for the 12-site hexagon. This must at least in part be due to the lack of extensivity, though it may also be related to the differently-shaped lattice. Unfortunately, the next larger hexagonal lattice contains 48 sites, which exceeds the capability of exact diagonalization to obtain the exact energy curve and there is no smaller rhombus that is both compatible with the 120°phase and possesses an $S=0$ ground state. As a result, we cannot generate exact results either which we can use to discriminate between effects due to lattice size and those due to lattice shape.
@@
Regardless of whether the lower accuracy seen in Fig. 10 is simply due to lack of extensivity or not, it is clear that projected AGP methods are inadequate in the thermodynamic limit. To address larger systems accurately, it is imperative to integrate correlated approaches, including J2CC, with the symmetry projection methods. This require further research. Alternatively, due to the superiority of symmetry-projected methods on small and medium-sized systems, it is also feasible to incorporate space group projection techniques with cluster mean field [39, 40] or density matrix embedding theory [41], where large systems are composed of fragments and SG-AGP may provide a good description of the fragments.

### v2

@@ I Introduction
Within the condensed matter physics community, spin model Hamiltonians have long served as a primary framework to study frustrated quantum magnetism [1, 2]. With a few exceptions [3], spin lattice models beyond one dimension are not exactly solvable, requiring the use of approximate numerical approaches instead. Analogously to the Hartree–Fock method used in the electronic structure theory of fermionic systems, spin wave functions based on a single spin configuration are insufficient for strongly correlated systems. Recent research has shown that the spin antisymmetrized geminal power (spin-AGP) method, which extends the AGP [4, 5] concept from systems of paired fermions to spin systems, may provide a good reference state for spin systems [6]. However, it is notable that, despite yielding some good results for one-dimensional (1D) systems, spin-AGP and correlated methods based on it are less accurate for lattices with more complicated geometry.
@@ II.2 Space group symmetry projection
$P_{LM}$ and $P_{PG}$ are the lattice momentum symmetry projection operator and point group symmetry projection operator, respectively. $T_{\mathbf{r}}$ is the translational operator with $\mathbf{q}$ being the lattice momentum projected onto. $R_{l}$ is a symmetry operator of the lattice point group and $\chi_{k}(l)$ is the character of the operator $R_{l}$ in irreducible representation $k$.
Note that space group operators in practice permute sites of the lattice. Therefore, space group operators acting on AGP produce another AGP whose geminal coefficients $\eta$ are permuted:
@@
For example, if the symmetry operator $G$ is the translation operator that translates site $i$ into $i+1$, and $\{\eta_{1},\eta_{2},\cdots,\eta_{M}\}$ is the geminal coefficient of the original AGP, we have $|\mathrm{AGP}(\eta_{G})\rangle=G|\mathrm{AGP}(\eta)\rangle=G|\eta_{1},\eta_{2},\cdots,\eta_{M}\rangle=|\eta_{2},\eta_{3},\cdots,\eta_{1}\rangle$.
@@
We note that the bra state can be simplified to the original AGP without permutation. Only the ket state need be explicitly projected, and therefore represented as a linear combination of non-orthogonal AGPs (LC-AGP). Since the number of space group symmetry operators, and hence the number of AGPs required for this LC-AGP, is typically $\mathcal{O}(M)$, where $M$ is the number of spin sites, the cost of space group symmetry-projected AGP is one factor of $M$ larger than that of the mean-field optimized AGP, making it $\mathcal{O}(M^{4})$. This is in contrast to a more general LC-AGP–style wave function, for which one needs to evaluate Hamiltonian overlaps between $\mathcal{O}(M^{2})$ pairs of AGPs, at an $\mathcal{O}(M^{5})$ computational cost.
@@ II.3 Complex Conjugate, Spin-Flip, and Time-Reversal Symmetry
where the complex conjugation operator $K$ is an anti-unitary operator and satisfies
@@
Because this set is closed, time-reversal symmetry is automatically projected once we apply the complex conjugation and spin-flip projectors.
@@ III.1 One-dimensional XXZ model
Figure 1: Energy of the 12-site 1D XXZ model with periodic boundary conditions by sBCS, AGP and space group projected AGP methods. SG-AGP with real $\eta$ is denoted as real-SG-AGP, while complex-LM-AGP and complex-SG-AGP represent LM-AGP and SG-AGP with complex $\eta$, respectively. These methods are compared with mean-field optimized AGP.
@@
The space group projected AGP is optimized with both real and complex geminal coefficient $\eta$. The energies of the 12-site XXZ chain with periodic boundary conditions ($M$=12) are shown in Fig. 1. The result of lattice momentum projected AGP (LM-AGP), which projects lattice momentum symmetry but not point group symmetry, with complex $\eta$ is also included. The results are compared with the mean-field optimized sBCS and AGP methods. Projecting $S^{z}$ to convert sBCS into AGP provides a noticeable but not large improvement to the total energy. However, the improvement upon incorporating space group projection as well is much larger. Furthermore, with complex $\eta$, the energy result is much superior to that of real $\eta$ at the cost of the complex conjugation symmetry. The errors of complex SG-AGP methods are shown in Fig. 2 and are compared with J2 coupled cluster (J2CC) [23], a correlated method based on a single AGP state. We note that the symmetry projection makes a huge difference in the energy results, and the complex-SG-AGP, as a symmetry-projected mean-field method, already produces a result comparable to correlated methods.
@@
Of course the energy is not the only quantity of interest. Correlation functions, in particular, are important and can help us understand the nature of magnetism. A single AGP is generally inadequate here, as it provides exact correlation functions at $\Delta=-1$ but predicts that the correlation function is independent of $\Delta$ for $|\Delta|\lesssim 1$.
Figure 3: $S^{z}$-$S^{z}$ correlation functions for the 12-site XXZ model with PBC, compared to those from SG-AGP with exact results. The lines are for exact results, and the circles are SG-AGP. Different colors represent different $\Delta$ values.
The $S^{z}-S^{z}$ spin-spin correlation is defined as
@@
where $i$ and $j$ are indices for sites. The results are shown in Fig. 3. It can be seen that the spin-spin correlation given by SG-AGP is quite close to exact, and as $\Delta$ goes from $-1$ to $1$, the system approaches a Néel structure.
Figure 4: AGP and SG-AGP overlap with full CI (FCI). The discontinuity in the AGP curve is because AGP has two solutions: one for $\Delta\lesssim 0.95$ and another for $\Delta\gtrsim 0.95$.
Figure 4 displays the overlap between the SG-AGP and the exact solution (FCI). This overlap is compared with that of AGP. While the AGP has relatively small overlap with the exact result except near $\Delta=-1$, the overlap of the SG-AGP and the exact solution is greater than 0.95 for all $\Delta$. This indicates that SG-AGP captures most of the exact wave function, and suggests that the remaining correlations atop SG-AGP, unlike those atop AGP, are small.
@@ III.2 Two-dimensional $\mathrm{J_{1}-J_{2}}$ model
where $\langle pq\rangle$ and $\langle\langle pq\rangle\rangle$ denote sites $p$ and $q$ being nearest neighbors and next-nearest neighbors respectively. We take $\mathrm{J_{1}}=1$, and the competition between nearest neighbors and next-nearest neighbors is only dependent on $\mathrm{J_{2}}$. The space group projected AGP methods are tested on the square and triangular lattices, both of which show rich physics and have spin-liquid [24] phases for intermediate $\mathrm{J_{2}}$.
@@ III.2.1 Square lattice
Figure 6: Energy of the $4\times 4$ square $\mathrm{J_{1}-J_{2}}$ model by space group projected AGP methods compared with mean-field optimized AGP and J2CC and sBCS-CCSD.
Figure 7: AGP, SG-AGP and SGKF-AGP overlap with full CI (FCI).
The energy of the $4\times 4$ square $\mathrm{J_{1}-J_{2}}$ model given by symmetry-projected AGP methods is shown in Fig. 6. We remind our readers that these energies are optimized with complex $\eta$ via the variation after projection approach. In addition to SG-AGP, we include the results from the combination of space group projection with complex conjugate and spin-flip projection, referred to as SGKF-AGP. Figure 7 demonstrates their overlap with the exact solution. Both SG-AGP and SGKF-AGP perform well in both the large and small $\mathrm{J_{2}}$ limits. In the most frustrated region around $\mathrm{J_{2}}\approx 0.5$, they maintain a respectable overlap with the exact solution and succeed in capturing a significant portion of the correlation energy. The energy results are compared with J2CC, which, despite the discontinuity in the highly frustrated region around $\Delta\sim 0.5$, provides some of our best results for correlation atop a single AGP. We also compare the results to spin BCS–based coupled cluster with single and double excitations (sBCS-CCSD), which is the standard approach for treating pair correlations, to offer readers a broader perspective on the accuracy of our methods. We see that with the symmetry projections, SG-AGP and SGKF-AGP can already surpass the performance of correlated methods based on a single AGP or spin BCS.
@@
We observe that when $\mathrm{J_{2}}=0$, there is a negative spin-spin correlation for the nearest neighbors, while the second-nearest neighbors exhibit a positive correlation. This is indicative of the checkerboard Néel phase. For $\mathrm{J_{2}}=1$, the second-nearest-neighbor correlation becomes negative while that of the nearest-neighbor tends toward 00, corresponding to the striped order. The projected AGP spin-spin correlation functions coincide well with the exact one for large and small $\mathrm{J_{2}}$. In the spin liquid phase with intermediate $\mathrm{J_{2}}$ values, the spin-spin correlation is qualitatively correct but shows some deviations for certain site distances.
Figure 8: $S^{z}$-$S^{z}$ correlation functions from various projected methods in the $4\times 4$ square $\mathrm{J_{1}-J_{2}}$ model with PBC, compared with exact results. Lines are for the exact result, crosses are for SG-AGP, and circles are for SGKF-AGP. Different colors represent different $\mathrm{J_{2}}$ values. The horizontal axis $j$ is arranged according to the distance between two spins where $j$ can be regarded as the $j$th-nearest-neighbor of site 00, as shown in the top right of the figure, where site 0 is colored black and 1—5 are the $j$th-nearest-neighbors, colored differently according to $j$.
@@ III.2.2 Triangular lattice
A typical 2D triangular lattice has a $D_{6}$ point group symmetry with 12 symmetry operations. For the same point symmetry to be obeyed, a finite-size lattice must be chosen to take the shape of either a rhombus or hexagon, as shown in Fig. 5. Lattice momentum symmetry involves translation along the 2 lattice vectors. The $D_{6}$ point group symmetry and the translation symmetry form the $p6m$ space group for the 2D triangular lattice and the SG-AGP for an $M$-site lattice is a linear combination of $12M$ AGP states.
@@
It is worth mentioning that in order to correctly capture the 120° phase, the lattice must have a multiple of 3 sites along each translation vector. We will test our methods on the 12-site and 36-site lattices, which satisfy this restriction, and we will also check its behavior on the 16-site lattice, which breaks this restriction as shown in the bottom left of Fig. 5: if different colors represent spins spaced by 120°, adjacent sites across the periodic boundary will showcase the same color, thereby breaching the 120° phase.
@@
Figure 9: Energy of the 12-site triangular $\mathrm{J_{1}-J_{2}}$ model by space group projected AGP methods compared with mean-field optimized AGP. The region of $0.260.2<\mathrm{J_{2}}<$ 0.26 is zoomed in at the bottom right of the figure.
@@
The 12-site lattice has a hexagonal shape, as shown in the top right of Fig. 5. The energy of space group projected AGP methods is shown in Fig. 9. Although a single mean-field optimized AGP provides only a qualitative insight, including additional symmetry projections brings the energy closer to the exact value, with the SG-AGP exhibiting an error of $10^{-2}$. Moreover, one of the exact eigenstates is obtained when the complex conjugate projection is included, which might be attributed to the limited size of the system. As shown in the bottom right of the figure, the SGK-AGP and the eigenstate feature the ground state for $\mathrm{J_{2}}<0.23$, after which a different eigenstate with a different spatial symmetry has a lower energy and SGK-AGP is no longer the ground state but still an eigenstate of the system. The spin-spin correlation of SG-AGP for this lattice is illustrated in Fig. 10, demonstrating that SG-AGP accurately captures the correlation.
@@
Figure 10: $S^{z}$-$S^{z}$ correlation functions of SG-AGP for the 12-site triangular $\mathrm{J_{1}-J_{2}}$ model compared with exact results. The lines are for the exact results, and the circles are for SG-AGP. Different colors represent different $\mathrm{J_{2}}$ values. The horizontal axis $j$ is arranged according to the distance between two spins where $j$ can be regarded as the $j$th-nearest-neighbor of site 00, as shown in the top right of the figure, where site 0 is colored black and 1—3 are the $j$th-nearest-neighbors, colored differently according to $j$.
@@
The space group projected AGP methods are also tested on the 16-site ($4\times 4$ rhombus) and 36-site ($6\times 6$ rhombus) triangular lattices shown in Fig. 11 and Fig. 12 respectively. In the 16-site lattice, where the lattice size is not a multiple of 3, due to the finite-size effect mentioned earlier, the energy curves of both the symmetry-projected AGP methods and the exact solution deviate from those observed in the 12-site and 36-site lattice models. Despite the absence of the 120° phase, the exact ground state of this lattice continues to exhibit the same space group symmetry as other system sizes. The space group projected AGP methods can accurately capture the energy, particularly SGKF-AGP, which has an error of approximately $10^{-1}$.
@@
However, an important limitation of projected AGP methods (and of AGP itself) is that they are not size extensive. In the limit of large lattices, the energy per site reduces to that of the spin-BCS. Figure 12 and Figure 13 shows energy and spin-spin correlation results for the 36-site rhombus lattice. We can see that the SG-AGP energy is significantly less accurate here than it is for the 12-site hexagon. The spin-spin correlation for the SGKF-AGP retains the correct sign, but it is not quantitatively accurate. The shortcomings for SG-AGP must at least in part be due to the lack of extensivity, though it may also be related to the differently-shaped lattice. Unfortunately, the next larger hexagonal lattice contains 48 sites and is too large for us to generate exact results against which we can compare, and there is no smaller rhombus that is both compatible with the 120° phase and possesses an $S^{z}=0$ ground state. As a result, we cannot generate exact data to discriminate between effects due to lattice size and those due to lattice shape.
@@
Regardless of whether the lower accuracy seen in Fig. 12 is simply due to lack of extensivity or not, it is clear that projected AGP methods are inadequate in the thermodynamic limit. To address larger systems accurately, it is imperative to integrate correlated approaches, including J2CC, with the symmetry projection methods. This requires further research. Alternatively, due to the superiority of symmetry-projected methods on small and medium-sized systems, it is also feasible to incorporate space group projection techniques with cluster mean field [39, 40] or density matrix embedding theory [41], where large systems are composed of fragments and SG-AGP may provide a good description of the fragments.
@@
Figure 13: $S^{z}$-$S^{z}$ correlation functions from various projected methods in the 36-site triangular $\mathrm{J_{1}-J_{2}}$ model with PBC, compared with exact results. Lines are for the exact result, crosses are for SG-AGP, and circles are for SGKF-AGP. Different colors represent different $\mathrm{J_{2}}$ values. The horizontal axis $j$ is arranged according to the distance between two spins where $j$ can be regarded as the $j$th-nearest-neighbor of site 00.